two bodies of unequal masses are thrown upwards with same velocity.if air exerts equal force on both bodies,will they reach same height or not?explain.?

Grade:12

3 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear mayank

Dear let force exert by air is F on both the bodies

do retardation = (F+mg)/m = g + F/m

clearly retardation on lighter body will  be more .so it will reach less hight as compair to heavier body.

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

arnab nandy
13 years ago

NO they won't

as far as gravitation's acc is considered they would have reached same height

now if these bodies experiences equal force

there different retarding acc of the air

M1 X A1 = M2 X A2

M1 NOT EQUAL TO M2

THEREFORE A1 NOT EQUAL TO A2

HENCE DIFFERENT HEIGHT THEY WILL REACH

SARATH NARAYAN PILLA
13 years ago

they wont reach same height.they hav 2 diff masses{say M and m}whose weight acts downwards i.e., Mgsinθ and mgsinθ.therefore whichever has more mass reaches less height.